OpenStudy (anonymous):

1) Given x^2-16=y^2 a) f''(x) in the simplest factored form b) f''(5)

Miracrown (miracrown):

f'' is the second derivative so we have to find the first derivative to start

Miracrown (miracrown):

\[x^2 - 16 = y^2\]

Miracrown (miracrown):

That is implicit ^. Would you like to take the derivative of both sides?

OpenStudy (anonymous):

yes

Miracrown (miracrown):

Ok, what will it be?

OpenStudy (anonymous):

d/dx(x^2-16)=d/dx(y^2)??

Miracrown (miracrown):

Yep, but for the last step ...

Miracrown (miracrown):

\[\frac{ d }{ dx } \space (x^2 - 16) \space = \frac{ d }{ dx } \space (y^2)\]

OpenStudy (anonymous):

is it x/y??

Miracrown (miracrown):

Would you like to try the next step? Our first goal is to get dy/dx (that will be f prime)

Miracrown (miracrown):

That's right! It is x/y... And that does make sense, this is a circle

Miracrown (miracrown):

\[f'(x) \space = \frac{ d }{ dx } \space = \frac{ x }{y }\]

Miracrown (miracrown):

so then we need to take the derivative once more to get f'' (second derivative)

OpenStudy (anonymous):

how do we do that?

Miracrown (miracrown):

You have two functions x and y that are divided, which means the quotient rule

OpenStudy (anonymous):

okay...

Miracrown (miracrown):

\[\frac{ d^2 y }{ dx^2 } \space = \frac{ y(1) - x \frac{ dy }{ dx } }{ y^2 }\]

OpenStudy (anonymous):

what do i do after?

Miracrown (miracrown):

ahhh, but you know dy/dx there, so let's substitute that also

Miracrown (miracrown):

\[= \frac{ y - x \space . \frac{ x }{ y } }{ y^2 }\]

Miracrown (miracrown):

and it wants it simplified, alrighty. Any ideas on what do first for that?

OpenStudy (anonymous):

get rid of the denominator?

Miracrown (miracrown):

that's what I tried first, too

OpenStudy (anonymous):

i'm not sure how to do that

Miracrown (miracrown):

You could get a common denominator in the numerator. To do that, we'd have to multiply the first fraction by y/y

OpenStudy (anonymous):

so first find the common denominator in the numerator

OpenStudy (anonymous):

so it would be y^2-x^2/y (1/y^2)

Miracrown (miracrown):

Yes

Miracrown (miracrown):

and then you could multiply that straight across

OpenStudy (anonymous):

it would be y^2-x^2/y^3

Miracrown (miracrown):

\[\frac{ y^2 - x^2 }{ y^3 }\]

Miracrown (miracrown):

Yep

OpenStudy (anonymous):

Yay! :)

Miracrown (miracrown):

:)

OpenStudy (anonymous):

how do i sub in the 5?

Miracrown (miracrown):

Right, because that is only the x = we'd have to solve for the y using the original equation

Miracrown (miracrown):

So \[5^2 - 16 = y^2\]

OpenStudy (anonymous):

so y is plus and minus 3?

Miracrown (miracrown):

yep, I got solutions y = 3, -3

OpenStudy (anonymous):

Quick question for the f''(x) i noticed in the equation of x^2-16=y^2 can be -16=y^2-x^2 so can we say that it's -16/y^3???

Miracrown (miracrown):

Oh, there the left hand side would have a derivative of 0 0 = 2y dy/dx - 2x

Miracrown (miracrown):

because of the constant

OpenStudy (anonymous):

so we can't do that?

Miracrown (miracrown):

Right, you can't because you have to take the derivative of both sides separately. o

Miracrown (miracrown):

Oh, but I see what you mean by plugging in! Hmm, sure... that's fine, actually.

Miracrown (miracrown):

You'll get the same answer either way \[\frac{ -16 }{ y^3 } \space \]

OpenStudy (anonymous):

so for this one, you can't sub in the 5?

Miracrown (miracrown):

oh, that's a good point. There is no x variable left after you simplify, so you would only need to sub y = 3

OpenStudy (anonymous):

So it would be 9-x^2/27?

Miracrown (miracrown):

Oh but you know what? y = root (x² - 16) right? So we could cube that and plug that in then if you wanted, you could sub x = 5 there using the same ideas you had before like with the -16 = y² - x2

Miracrown (miracrown):

it would be, yep, but you could keep the -16 in the numerator, too

Miracrown (miracrown):

All of these are correct actually.

Miracrown (miracrown):

Then plugging x = 5 would work, agree

OpenStudy (anonymous):

oohhh i see

OpenStudy (anonymous):

so for f''(5) answers are plus and minus 16/27??

OpenStudy (anonymous):

@Miracrown can you check if it's right? And thank you for helping me out! Also thanks for answering my dummy questions. :)

Miracrown (miracrown):

That looks good to me. :)

OpenStudy (anonymous):

:D